Case: Cost Structures for Global Shippers Inc.

Management from Global Shippers Inc, an international shipping business, is in the process of assessing the choice between two different cost structures for the business. Option A has relatively higher variable costs per unit shipped but lower annual fixed costs, while Option B has the opposite—relatively lower variable costs in its cost structure but higher fixed costs. Assume that delivery selling prices per unit are constant. The table below contains critical information in making the decision:
Cost Information Option A Option B
Delivery price (revenue) per shipment $100 $100
Variable cost per shipment delivered $85 $60
Contribution Margin per unit $15 $40
Fixed costs (annual) $1,450,000 $4,700,000

Management wants you to write a professional report, answering the following questions:
Questions
1) What is the break-even point, in terms of volume (i.e., number of shipments per year), for Option A? Option B?
(2) How many shipments would have to be made under Option A to produce operating income of $43,000 for an annual period?
(3) How many shipments per month would have to be made under Option A to produce an annual operating margin equal to 11% of sales revenue?
(4) How many shipments are required under Option B to produce net income of $213,000 per year, given a corporate tax rate of 25%?
(5) Assume that for the coming year total fixed costs are expected to decrease by 8% for each of the two options. What is the new break-even point, in terms of number of shipments, for each option? By what percentage did the break-even point change for each case? How do these figures compare to the percentage increase in budgeted fixed costs?
(6) Assume an average income-tax rate of 35%. What volume (number of shipments) would be needed to generate net income of 8% of revenue for each option?
(7) Which option do you think is the more profitable one for this business? Explain.
(8) Which option do you consider to be more risky to the business? Explain (calculate degree of operating leverage to help answer this question).
Answer #1

Part 1

Breakeven point = fixed costs / contribution margin per unit

Option A = 1450000/15 = 96667

Option B = 4700000/40 = 117500

Part 2

Desired shipments = (fixed costs + targeted operating income)/ contribution margin per unit

= (1450000+43000)/15 = 99533

Part 3

Profit = SX – VCX – FC

0.11X = 100X-85X-1450000

14.89X = 1450000

X = 97381

Part 4

Operating income = net income / (1-t) = 213000/(1-25%) = $284000

Desired shipments = (fixed costs + targeted operating income)/ contribution margin per unit = (4700000+284000)/40 = 124600

Part 5

Option A

New fixed cost = 1450000*92% = 1334000

Breakeven point = 1334000/15 = 88933

Decrease in shipments = (96667-88933)/96667 = 8.00%

Option B

New fixed cost = 4700000*92% = 4324000

Breakeven point = 4324000/40 = 108100

Decrease in shipments = (117500-108100)/117500 = 8.00%

With percentage increase in budgeted fixed costs, the number of shipments required to breakeven will also increase by the same percentage of increase in budgeted fixed costs. In short increase or decrease in percentage of change in breakeven point varies directly to the percentage increase or decrease in fixed costs.

Part 6

Option A

Profit = SX – VCX – FC

(8/65)X = 100X-85X-1450000

8X=975X-94250000

967X = 94250000

X = 97466

Option B

Profit = SX – VCX – FC

(8/65)X = 100X-60X-4700000

8X = 2600X -30550000

2592 X = 30550000

X = 11786

Part 7

Option B. The option with higher contribution margin is more profitable for the business.

Part 8

degree of operating leverage = contribution margin / net operating income

option A = (97466*15)/((97466*15)-1450000) = 121.93

option B = (11786*40)/((11786*40)-4700000) = 327.39

option B is more risky because it has higher proportion of fixed asset resulting into the higher DOL